OpenStudy (anonymous):

Explain, in complete sentences, how you would factor \(32x^6 – 162\) Provide the factors. I am seriously confused.

OpenStudy (anonymous):

At the least you can post the question correctly.

OpenStudy (anonymous):

Doh

OpenStudy (anonymous):

There.

OpenStudy (anonymous):

\(32x^6 – 162 = 2(16x^6-81) =2 \left(4 x^3-9\right) \left(4 x^3+9\right) \)

OpenStudy (anonymous):

Difference of square format?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Thank you! :)